# TransBox: $\mathcal{EL}^{++}$-closed Ontology Embedding

### v1

@@ Abstract.
OWL (Web Ontology Language) ontologies, which are able to represent both relational and type facts as standard knowledge graphs and complex domain knowledge in Description Logic (DL) axioms, are widely adopted in domains such as healthcare and bioinformatics. Inspired by the success of knowledge graph embeddings, embedding OWL ontologies has gained significant attention in recent years. Current methods primarily focus on learning embeddings for atomic concepts and roles, enabling the evaluation based on normalized axioms through specially designed score functions. However, they often neglect the embedding of complex concepts, making it difficult to infer with more intricate axioms. This limitation reduces their effectiveness in advanced reasoning tasks, such as Ontology Learning and ontology-mediated Query Answering. In this paper, we propose $\mathcal{EL}^{++}$-closed ontology embeddings which are able to represent any logical expressions in DL $\mathcal{EL}^{++}$ via composition. Furthermore, we develop TransBox, an effective $\mathcal{EL}^{++}$-closed ontology embedding method that can handle many-to-one, one-to-many and many-to-many relations. Our extensive experiments demonstrate that TransBox often achieves state-of-the-art performance across various real-world datasets for predicting complex axioms.
@@ 2. Preliminaries and Related Work
2.1. Ontologies
Ontologies use sets of statements (axioms) about concepts (unary predicates) and roles (binary predicates) for knowledge representation and reasoning. We focus on $\mathcal{EL}^{++}$-ontologies which keep a good balance between expressivity and reasoning efficiency with wide application (Baader and Gil, 2024). Let $\mathsf{N_{C}}=\{A,B,\ldots\}$, $\mathsf{N_{R}}=\{r,t,\ldots\}$, and $\mathsf{N_{I}}$=$\{a,b,\ldots\}$ be pair-wise disjoint sets of concept names (also called atomic concepts) and role names, and individual names, respectively. $\mathcal{EL}^{++}$-concepts are recursively defined from atomic concepts, roles and individuals as
(1) $\top~{}|~{}\bot~{}|~{}A~{}|~{}C\sqcap D~{}|~{}\exists r.C~{}|\left\{a\right\}$
and an $\mathcal{EL}^{++}$-ontology is a finite set of TBox axioms of the form
$C\sqsubseteq D,\quad r\sqsubseteq t,\quad r_{1}\circ r_{2}\sqsubseteq t$
and ABox axioms $A(a)$ or $r(a,b)$. Note $C,D$ are (possibly complex) $\mathcal{EL}^{++}$-concepts, $A$ is a concept name, $r_{1},r_{2},t$ are role names, and $a,b$ are individual names. The following is an example of axioms of a toy family ontology (see Figure 5 for the complete version).
@@ 3. $\mathcal{EL}^{++}$-Closed Embeddings
The $\mathcal{EL}^{++}$-closed embeddings refer to embedding methods where the assigned spaces $\mathcal{E}_{C}$ (for concepts) and $\mathcal{E}_{R}$ (for roles) are closed under the $\mathcal{EL}^{++}$-semantics defined by Equations 2, 16, and 4. The formal definition is provided in Definition 3.2, where each of the three conditions directly corresponds to these equations.
@@ 4.3. Semantic Enhancement
$Box(B)$$Box(\exists^{\text{all}}_{r}B)$$Box(r)$
Figure 3. Illustration of $Box(\exists^{\text{all}}_{r}B)$
(a) ELBE
(b) TransBox (w/o Enhancement)
(c) TransBox
@@
Figure 5. Family Ontology
@@ 5. Soundness
Let $\mathcal{O}$ be an $\mathcal{EL}^{++}$ ontology. We can extend any TransBox embedding of $\mathcal{O}$ to an interpretation $\mathcal{I}_{g}$ following the $\mathcal{EL}^{++}$ semantics introduced in Section 2.1, where the atomic concepts in $\mathcal{O}$ are represented by $Box(A)$, and the roles in $\mathcal{O}$ by $E_{Box(r)}$, as specified in Equation (6).
@@ 6.1. Proof of Concept: Family Ontology
To evaluate and demonstrate the expressiveness of the TransBox embedding method, we use a simple family ontology and compare the embedding results with ELBE (Peng et al., [n. d.]). For better illustration, we train all embeddings in a 2-dimensional space, and add visualization loss below into the final loss functions to avoid overly small boxes as in Box2EL (Jackermeier et al., 2024).
$\mathcal{L}_{V}=\frac{1}{n|\textsf{N}_{C}|}\sum_{A\in\textsf{N}_{C}}\sum_{i=1}^{n}\max\{0,0.2-o(Box(A))_{i}\}.$
In this simple experiment, we set the margin to $\gamma=0$, regularization factor $\lambda=0$, and omit negative sampling. The resulting embeddings are shown in Figure 4.
@@
In contrast, TransBox embeddings (Figures 4(b) and 4(c)) resolve this issue by utilizing multi-transition representations within $Box(r)$. This approach allows TransBox to correctly learn distinct embeddings for each concept. Additionally, applying the semantic enhancements described in Section 4.3 further improves the embeddings of roles. Specifically, in Figure 4(c), we observe that the embedding $Box(hasParent)\approx-Box(hasChild)$ well captures the fact that the role hasParent is the inverse of hasChild.
@@ Results
The overall result is summarized in Table 2, and the result of each subtask is presented in Tables 3, 4, 5 and 6. Note that for the task of predicting complex axioms, we can only make a comparison with $\mathcal{EL}^{++}$-closed methods BoxEL and ELBE.
@@
Table 7. Overall comparison of $\mathcal{EL}$-closed method on normalized axioms
@@ 6.2.2. Normalized Axioms Prediction
The overall results for the four types of normalized axioms ($A\sqsubseteq B$, $A\sqsubseteq B\sqsubseteq B^{\prime}$, $A\sqsubseteq\exists r.B$, and $\exists r.B\sqsubseteq A$) are presented in Table 7. While TransBox achieved the best overall performance among $\mathcal{EL}^{++}$ methods, it is less competitive compared to $\mathcal{EL}^{++}$ methods for predicting normalized axioms. This is expected, as $\mathcal{EL}^{++}$ methods, unconstrained by extensional capabilities, are often better suited to capture normalized axioms due to their specific design. For instance, Box2EL’s use of bump vectors provides enhanced capacity for embedding concepts and roles to fit axioms. However, this approach is restricted to atomic concepts for which bump vectors are explicitly defined. Furthermore, it introduces greater model complexity for storing these bump vectors compared to TransBox, with a total complexity of $O(n(3N_{C}+4N_{R}+2N_{I}))$ versus $O(n(2N_{C}+2N_{R}+N_{I}))$.
@@ 6.2.4. Case Study
In the following query of the form $C\sqsubseteq?A$ over GALEN, our model correctly identified the correct answer EndometriosisLesion with a score of -0.32, ranking 3rd. Note the score is the negative value of the distance. It effectively distinguished EndometriosisLesion from other similar conditions such as EndometrialHypoplasia (-0.65), EndometrialRegeneration (-0.69), and EndometrialNeoplasia (-0.64).
@@
This example also underscores the significance of handling complex axioms in practical, real-world applications. In the original, normalized GALEN ontology, EndometriosisLesion appears on the right-hand side of two normalized axioms:
[…]
While these normalized axioms are informative, the more complex axiom used in the query provides richer context and deeper insights that may not be captured in simpler forms. Additionally, this complexity makes the axiom more intuitive and actionable for medical professionals, aiding in more accurate diagnosis and decision-making.
@@ Appendix A Proofs
Appendix B Box2EL treat roles as transitive
@@ Appendix D Generation of complex axiom
In detail, for each ontology $\mathcal{O}$, we create a signature $\Sigma$ consisting of 1,000 randomly selected atomic concepts. Then, we perform forgetting over $\mathcal{O}$ and $\Sigma$, resulting in a forgotten ontology $\mathcal{O}_{\Sigma}$. Finally, we select all axioms in $\mathcal{O}_{\Sigma}$ with a length [^4] We define the length of an axiom as the number of atomic concepts and roles it contains between 4 and 10. We repeat the progress multiple times if the generated axioms are less than 1000. The exact progress has been shown in Algorithm 1.

### v2

@@ Abstract.
OWL (Web Ontology Language) ontologies, which are able to represent both relational and type facts as standard knowledge graphs and complex domain knowledge in Description Logic (DL) axioms, are widely adopted in domains such as healthcare and bioinformatics. Inspired by the success of knowledge graph embeddings, embedding OWL ontologies has gained significant attention in recent years. Current methods primarily focus on learning embeddings for atomic concepts and roles, enabling the evaluation based on normalized axioms through specially designed score functions. However, they often neglect the embedding of complex concepts, making it difficult to infer with more intricate axioms. This limitation reduces their effectiveness in advanced reasoning tasks, such as Ontology Learning and ontology-mediated Query Answering. In this paper, we propose $\mathcal{EL}^{++}$-closed ontology embeddings which are able to represent any logical expressions in DL $\mathcal{EL}^{++}$ via composition. Furthermore, we develop TransBox, an effective $\mathcal{EL}^{++}$-closed ontology embedding method that can handle many-to-one, one-to-many and many-to-many relations. Our extensive experiments demonstrate that TransBox often achieves state-of-the-art performance across various real-world datasets for predicting complex axioms [^1] Code and data are available at the https://github.com/HuiYang1997/TransBox.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714672isbn: 979-8-4007-1274-6/25/04ccs: Information systems Web Ontology Language (OWL)ccs: Theory of computation Description logicsccs: Computing methodologies Ontology engineeringccs: Computing methodologies Learning latent representationsccs: Computing methodologies Statistical relational learning
@@ 2. Preliminaries and Related Work
Ontologies
Ontologies use sets of statements (axioms) about concepts (unary predicates) and roles (binary predicates) for knowledge representation and reasoning. We focus on $\mathcal{EL}^{++}$-ontologies which keep a good balance between expressivity and reasoning efficiency with wide application (Baader and Gil, 2024). Let $\mathsf{N_{C}}=\{A,B,\ldots\}$, $\mathsf{N_{R}}=\{r,t,\ldots\}$, and $\mathsf{N_{I}}$=$\{a,b,\ldots\}$ be pair-wise disjoint sets of concept names (also called atomic concepts) and role names, and individual names, respectively. $\mathcal{EL}^{++}$-concepts are recursively defined from atomic concepts, roles, and individuals as $\top~{}|~{}\bot~{}|~{}A~{}|~{}C\sqcap D~{}|~{}\exists r.C~{}|\left\{a\right\}$ and an $\mathcal{EL}^{++}$-ontology is a finite set of TBox axioms of the form $C\sqsubseteq D,r\sqsubseteq t,r_{1}\circ r_{2}\sqsubseteq t$ and ABox axioms $A(a)$ or $r(a,b)$. Note $C,D$ are (possibly complex) $\mathcal{EL}^{++}$-concepts, $A$ is a concept name, $r_{1},r_{2},t$ are role names, and $a,b$ are individual names. The following is an example of axioms of a toy family ontology (cf. Figure 4(a)).
@@ 3. $\mathcal{EL}^{++}$-Closed Embeddings
The $\mathcal{EL}^{++}$-closed embeddings refer to embedding methods where the assigned spaces $\mathcal{E}_{C}$ (for concepts) and $\mathcal{E}_{R}$ (for roles) are closed under the $\mathcal{EL}^{++}$-semantics defined by Equations 1, 2, and 3. The formal definition is provided in Definition 3.2, where each of the three conditions directly corresponds to these equations.
@@ 4. Method: TransBox
$Box(B)$$Box(\exists r.B)$$Box(r)$
Figure 2. Illustration of $Box(\exists r.B)$
$Box(B)$$Box(\exists^{\text{all}}_{r}B)$$Box(r)$
Figure 3. Illustration of $Box(\exists^{\text{all}}_{r}B)$
@@ 4.3. Semantic Enhancement
[…]
(a) Family Ontology
(b) ELBE
(c) TransBox (w/o Enhancement)
(d) TransBox
@@ 5. Soundness
Let $\mathcal{O}$ be an $\mathcal{EL}^{++}$ ontology. We can extend any TransBox embedding of $\mathcal{O}$ to an interpretation $\mathcal{I}_{g}$ following the $\mathcal{EL}^{++}$ semantics introduced in Section 2, where the atomic concepts in $\mathcal{O}$ are represented by $Box(A)$, and the roles in $\mathcal{O}$ by $E_{Box(r)}$, as specified in Equation (5).
@@ 6.1. Proof of Concept: Family Ontology
To evaluate and demonstrate the expressiveness of the TransBox embedding method, we use a simple family ontology and compare the embedding results with ELBE (Peng et al., [n. d.]). For better illustration, we train all embeddings in a 2-dimensional space, and add visualization loss below into the final loss functions to avoid overly small boxes as in Box2EL (Jackermeier et al., 2024) (see $\mathcal{L}_{V}$ in Section 4.1). In this simple experiment, we set the margin to $\gamma=0$, regularization factor $\lambda=0$, and omit negative sampling. The resulting embeddings are shown in Figure 4.
@@
In contrast, TransBox embeddings (Figures 4(c) and 4(d)) resolve this issue by utilizing multi-transition representations within $Box(r)$. This approach allows TransBox to correctly learn distinct embeddings for each concept. Additionally, applying the semantic enhancements described in Section 4.3 further improves the embeddings of roles. Specifically, in Figure 4(d), we observe that the embedding $Box(hasParent)\approx-Box(hasChild)$ well captures the fact that the role hasParent is the inverse of hasChild.
@@ Results
The overall result is summarized in Table 2, and the result of each subtask is presented in Tables 3, 4, 5 and 6. Note that for the task of predicting complex axioms, we can only make a comparison with $\mathcal{EL}^{++}$-closed methods BoxEL and ELBE. Box2EL was excluded because it cannot handle complex axioms. Specifically, it fails to derive the embedding of $\exists r.B$ from the embeddings of $r$ and $D$.
@@
Table 7. Overall comparison on normalized axioms
@@ 6.2.2. Normalized Axioms Prediction
The overall results for the four types of normalized axioms ($A\sqsubseteq B$, $A\sqsubseteq B\sqsubseteq B^{\prime}$, $A\sqsubseteq\exists r.B$, and $\exists r.B\sqsubseteq A$) are presented in Table 7. While TransBox achieved the best overall performance among $\mathcal{EL}^{++}$-closed methods, it is less competitive compared to non $\mathcal{EL}^{++}$-closed methods for predicting normalized axioms. This is expected, as non $\mathcal{EL}^{++}$-closed methods, unconstrained by extensional capabilities, are often better suited to capture normalized axioms due to their specific design. For instance, Box2EL’s use of bump vectors provides enhanced capacity for embedding concepts and roles to fit axioms. However, this approach is restricted to atomic concepts for which bump vectors are explicitly defined. Furthermore, it introduces greater model complexity for storing these bump vectors compared to TransBox, with a total complexity of $O(n(3N_{C}+4N_{R}+2N_{I}))$ versus $O(n(2N_{C}+2N_{R}+N_{I}))$.
@@
Due to space constraints, the comparison with ball-based methods and the performance on each type of axiom are deferred to Appendix G. Briefly, ball-based methods achieved slightly lower but comparable performance to Box2EL and generally outperformed TransBox in predicting normalized axioms. This may be attributed to their lower parameter count for embeddings of the same dimension, making them more stable and suitable for training.
@@ 7. Conclusion and Future Work
ACKNOWLEDGEMENTS
This work is funded by EPSRC projects OntoEm (EP/Y017706/1) and ConCur (EP/V050869/1).
@@ Appendix A Proofs
A.6. Theorem 5.1
Proof.
Appendix B Box2EL reat roles as transitive
@@ Appendix D Generation of complex axiom
In detail, for each ontology $\mathcal{O}$, we create a signature $\Sigma$ consisting of 1,000 randomly selected atomic concepts. Then, we perform forgetting over $\mathcal{O}$ and $\Sigma$, resulting in a forgotten ontology $\mathcal{O}_{\Sigma}$. Finally, we select all axioms in $\mathcal{O}_{\Sigma}$ with a length between 4 and 10. Here, the length of an axiom is the number of atomic concepts and roles it contains. We repeat the progress multiple times if the generated axioms are less than 1000. The exact progress has been shown in Algorithm 1.